OpenStudy (anonymous):

Will give medal! 5+4x+2 ge x+8-1

OpenStudy (anonymous):

what is \(5+4x+2=\) ?

OpenStudy (anonymous):

which really just means, what is \(5+2\) ?

OpenStudy (anonymous):

3 hah aha

OpenStudy (anonymous):

lol no

OpenStudy (anonymous):

OH 7 srry just tired

OpenStudy (anonymous):

the point is , you need to start with \[4x+7\geq x+7\]

OpenStudy (anonymous):

subtract \(x\) from both sides to get \[3x+7\geq 7\] subtract 7 and get \[3x\geq 0\] then divide by 3 to get \[x\geq 0\]